# A Comprehensive Survey of Synthetic Tabular Data Generation

### v2

@@ III-B2 Score-based Method
Score-based method learns a score function, which represents the gradient of the logarithmic density of the data distribution. Unlike traditional diffusion models that rely on variational bounds and discrete-time Markov chains, score-based models use the gradient of the log density score function to define and reverse a continuous-time diffusion trajectory. Let $x(t)$ be a continuous-time variable that follows forward SDE [102, 103]:
@@
$dx=\left[f(x,t)-g(t)^{2}\nabla_{x}\log p_{t}(x)\right]dt+g(t)\,dw,$ (11)
where $\nabla_{x}\log p_{t}(x)$ is the score function of the data distribution at time $t$, and $dw$ is the reverse-time Wiener process. Since $\log p_{t}(x)$ is unknown, it is estimated using a score network $s_{\theta}(x,t)\approx\nabla_{x}\log p_{t}(x)$, trained using denoising score matching:
$\begin{split}\mathcal{L}(\theta)=&\;\mathbb{E}_{t\sim\mathcal{U}(0,T)}\lambda(t)\,\mathbb{E}_{x_{0},x(t)}\big{[}\|s_{\theta}(x(t),t)\\ &\;-\nabla_{x}\log p_{t}(x(t)\mid x_{0})\|_{2}^{2}\Big{]}\end{split}$ (12)
where $\lambda(t)$ is a weighting function, and $x_{0}\sim p_{\text{data}}$, and $x(t)$ is the noisy version of $x_{0}$. Once trained, the model generates samples by solving the reverse SDE (e.g., using numerical solvers like Euler-Maruyama [104] or Predictor-Corrector schemes [105]), starting from Gaussian noise $x(T)\sim\mathcal{N}(0,I)$ and integrating backward in time. These models have gained attention for their flexibility and effectiveness in capturing complex data distributions, making them well-suited for generating high-fidelity synthetic data across various domains. SOS [65] is the first score-based tabular data oversampling method. SOS [65] has redesigned the network to process tabular data. And through fine-tuning, the oversampling quality has been further improved. MissDiff [66] introduces a unified diffusion-based framework that incorporates a masked regression loss for denoising score matching during the training phase. This approach enables the model to effectively learn from data with missing values, addressing various missing data mechanisms and improving the quality of synthetic data in the presence of incomplete information. To effectively process both categorical and numerical variables simultaneously, CSDI-T [67] explores three techniques: one-hot encoding, analog bit encoding, and feature tagging. STaSy [68] leverages self-paced learning techniques and fine-tuning strategies to enhance the stability of denoising score-matching training, leading to improved sampling quality and greater diversity in synthetic tabular data. Forest-VP/Forest-Flow [69] employs XGBoost [106], leveraging its strength in tabular data prediction and classification. Importantly, XGBoost [106] can naturally handle missing data by learning the best split, allowing the model to be trained directly on incomplete datasets—an advantage over most generative models that require complete data. TABSYN [70] synthesizes tabular data by leveraging a diffusion model within a variational autoencoder (VAE) crafted latent space.
@@ III-C LLM-based Methods
Generating high-quality and realistic tabular data, particularly in specialized domains such as finance and healthcare, often requires domain-specific knowledge. For instance, generating synthetic heart disease records necessitates an understanding that the probability of heart disease increases with age, alongside general knowledge, such as the fact that a person’s age cannot be negative. However, existing generative models typically focus on learning the statistical distribution of raw tabular data without incorporating such knowledge. As a result, they may generate data that aligns with the learned distribution but violates logical or domain-specific constraints. To address this challenge, recent research explores leveraging the knowledge embedded in extensively pre-trained LLMs for synthetic tabular data generation, an overview of LLM-based models for synthetic tabular data generation is summarized in Table III. Depending on whether the LLM is fine-tuned, these methods can be broadly categorized into Prompt-based methods and Fine-Tuning methods as shown in Figure 4.
@@ III-C1 Prompt-based Methods
In-context learning (ICL) offers an innovative approach by enabling language models to perform tasks based solely on input-output examples, without parameter updates or fine-tuning. This capability, first observed in GPT-3 and subsequent LLMs [107], allows models to generalize to new tasks by embedding examples directly into the input prompt, which is often referred to as an "emergent ability" [108], this phenomenon has spurred significant research into understanding and enhancing ICL in models exceeding 100 billion parameters. Leveraging this capability, existing methods serialize the existing tabular dataset $D_{e}$ into a format compatible with LLM’s input. By carefully designing generation prompts, these methods guide the LLM to synthesize synthetic tabular dataset $D_{s}$. EPIC [71] leverages balanced, grouped data samples and consistent formatting with unique variable mapping to guide LLMs in generating accurate synthetic data across all classes, even for imbalanced datasets. CLLM [16] leverages the prior knowledge of LLMs for data augmentation in the low-data regime, and introduces a principled curation mechanism, leveraging learning dynamics, coupled with confidence and uncertainty metrics, to obtain a high-quality dataset. LITO [72] proposes an oversampling framework for tabular data to guide the abilities of generative language models. By leveraging its conditional sampling capabilities, LITO [72] synthesizes minority samples by progressively masking the important features of the majority class samples and imputing them towards the minority distribution.
@@ III-C2 Fine-Tuning Methods
Directly prompting LLMs for synthetic tabular data generation may lead to inaccurate or misleading outputs that deviate from user instructions. To mitigate this issue, recent research has explored fine-tuning LLMs on tabular data, enabling them to better understand the structure, constraints, and relationships within tabular datasets. Fine-tuning helps improve the accuracy and coherence of synthetic tables, ensuring that the output aligns more closely with real-world data distributions and domain-specific knowledge. GReaT [74] fine-tune LLM for sampling synthesized but highly realistic tabular data. In addition, GReaT can model the distribution of tabular data by adjusting any subset of features. TabMT [75] is a Bert-based mask transformer design used to generate synthetic tabular data, effectively addressing the unique challenges posed by heterogeneous data fields and capable of handling lost data. TAPTAP [76] utilizes tabular data pre-training to enhance the model’s tabular prediction ability. After pre-training a large amount of real-world tabular data, TAPTAP [76] can generate high-quality tabular data to support various applications, including privacy protection, low resource management, missing value imputation, and imbalanced classification. HARMONIC [77] constructs an instruction fine-tuning dataset based on the idea of the k-nearest neighbors algorithm to inspire LLMs to discover inter-row relationships. By fine-tuning, LLMs are trained to remember the format and connections of the data rather than the data itself, which reduces the risk of privacy leakage. AIGT [78] based on prompt enhancement, which utilizes metadata information, such as tabular descriptions and schemas, as prompts to generate a synthetic tabular dataset $D_{s}$ and propose long token partitioning algorithms that enable AIGT to model tables of any scale. DP-LLMTGen [17] for differentially private tabular data synthesis that leverages pre-trained large language models (LLMs). DP-LLMTGen models sensitive datasets using a two-stage fine-tuning procedure with a novel loss function specifically designed for tabular data. Subsequently, it generates synthetic data through sampling the fine-tuned LLMs. P-TA [82] proposes the use of Proximal Policy Optimization (PPO [109]) to apply GANs to guide LLMs to refine the probability distribution of tabular features and thereby generate more realistic data.
@@ IV Post Processing
The generative model introduced in Section 3 is theoretically capable of generating an infinite amount of tabular data. However, these synthetic samples often contain a number of unrealistic, incorrect, or even toxic instances. This issue arises from two main factors. First, the distribution of tabular data is highly complex, making its generation a traditionally challenging task. Second, non-LLM-based methods lack common sense and professional knowledge, leading to the generation of data that conforms to the statistical distribution but deviates from human knowledge. In LLM-based tabular generation methods, the "hallucination problem" exacerbates this issue, as LLMs inevitably produce flawed samples with incorrect labels. To address these challenges, several post-processing techniques for synthetic tabular data have been proposed. These methods can be broadly categorized into two groups: sample enhancement and label enhancement, as shown in Figure 5.
@@ IV-A Sample Enhancement
Sample enhancement methods aim to improve the quality and rationality of synthetic tabular data by modifying feature values in the synthetic samples or filtering the unreasonable samples. These methods aim to transform originally unreasonable or invalid samples into more realistic ones. They can be broadly categorized into two main approaches: sample filtering and sample correction.
@@ IV-A1 Sample Filtering
For sample filtering methods, the key step is to design appropriate criteria based on learning dynamics [110, 111, 112], such as confidence scores, impact functions, and generation capability, and filter samples under the criteria. CLLM [16] uses learning dynamics to provide significant signals about sample value for generative models. In order to provide intuition, samples with variable predictions may be considered fuzzy, or other samples may never learn correctly and may damage the model, and then these dynamics are analyzed by calculating two key metrics: confidence and arbitrary (data) uncertainty. These indicators form the basis for analyzing synthetic samples. LITO [72] employs self-authentication with rejection sampling, leveraging the intrinsic properties of language models. Specifically, even when a label is absent, the language model can infer and generate a label for a given sample. By inputting the labels of synthetic samples, the model itself can validate their correctness, effectively filtering out inaccurate or inconsistent outputs. This approach enhances the reliability of synthetic tabular data without requiring external validation mechanisms.
@@ IV-A2 Sample Correction
Sample correction incorporates neuro-symbolic principles into the training process of tabular data generation models by introducing human knowledge in the form of constraints or rules. Such, LL [113] addresses the issue of aligning synthetic test samples with background knowledge by adding an additional layer to the generative model. This layer restricts the output space to comply with a set of linear inequalities, ensuring consistency between synthetic data and available domain knowledge. However, this approach is limited to cases where the knowledge can be accurately expressed using linear inequalities, whose representational capacity is inherently constrained. To overcome this limitation, DRL [114] extends the approach by representing background knowledge using Quantifier-Free Linear Real Arithmetic (QFLRA) formulas, allowing the model to enforce constraints in non-convex and even disconnected spaces.
@@ IV-B Label Enhancement
Label enhancement methods aim to correct potential annotation errors in synthetic samples. The most straightforward approach is manual re-annotation of mislabeled data. Wang et al. (2021) [115] propose actively selecting samples with the lowest confidence scores for human re-labeling, while Pangakis et al. (2023) [116] and Liu et al. (2022a) [117] further emphasize the importance of manual review in ensuring data quality. However, despite its effectiveness, manual re-labeling incurs significant costs and is often impractical for large-scale deployment. To address these limitations, most synthetic tabular data generation methods adopt the proxy model approach, which leverages automated models to refine and correct labels. TAPTAP [76] and AIGT [78] use generative models to generate table features and then use the backbone discriminative model trained on the original data to generate pseudo labels for the synthetic data. Pred-LLM [79] generates table features through pre-training LLM, while further prompting LLM to generate its labels.
@@ V Evaluation
Kaggel Churn CH 10000 11(7/4) None Binclass Finance
@@
In the section, we categorize the evaluation of synthetic tabular data generation into two dimensions: Data Availability and Privacy. Data availability refers to the effectiveness of the synthetic data in supporting downstream machine learning tasks, including learning efficiency, fidelity to the original data distribution, and alignment with human domain knowledge. Privacy, on the other hand, measures the extent to which the synthetic data avoids disclosing sensitive information that could potentially be used to re-identify user information from the original dataset.
@@ V-B1 ML Efficiency
Machine Learning (ML) utility is a key criterion for evaluating the effectiveness of synthetic tabular data, as it reflects the extent to which the synthetic data can support downstream learning tasks. One widely adopted evaluation protocol for this purpose is the Training on Synthetic and Testing on Real (TSTR) scheme [118]. In this framework, the original dataset is first divided into a training set and a test set. A generative model is then trained on the original training data and used to produce a synthetic tabular dataset $D_{s}$ of the same size. Subsequently, two machine learning models are trained separately: one on the real training data and the other on the synthetic data. Both models are then evaluated using the same real test set. The core assumption behind this scheme is that if the synthetic data accurately captures the underlying patterns and distributions of the original data, then a model trained on the synthetic data should generalize well to real-world data. Therefore, if the performance of the model trained on synthetic data is comparable to that of the model trained on real data (e.g., in terms of accuracy, F1 score, AUC, etc.), the synthetic tabular dataset $D_{s}$ is deemed to possess high machine learning utility.
@@ V-B2 Fidelity
Fidelity measures how well synthetic data preserves the statistical properties of the original dataset. Although recent studies [119, 120] suggest that high fidelity does not always imply strong downstream task performance, it remains a crucial aspect of evaluating synthetic data quality.
@@
  • Column-wise: Compares individual features between real and synthetic data. Common metrics include Wasserstein distance [94] and Kolmogorov-Sirmov Test (KST) [121] for continuous features, and Jensen-Shannon Divergence (JSD) [122] or Total Variation Distance (TVD) [123] for categorical ones.
  • Pair-wise: Assesses how well dependencies between feature pairs are preserved using metrics such as Pearson Correlation Coefficient Matrices (DPCM) (for continuous features), The difference between Contingency Similarity Matrices among categorical features (DCSM) (for categorical features), and The difference between Contingency Similarity Matrices (for continuous and categorical features).

### v3

@@ III-B2 Score-based Method
Score-based method learns a score function, which represents the gradient of the logarithmic density of the data distribution. Unlike traditional diffusion models that rely on variational bounds and discrete-time Markov chains, score-based models use the gradient of the log density score function to define and reverse a continuous-time diffusion trajectory. Let $x$ be a continuous-time feature that follows forward SDE [106, 107]:
@@
$dx=\left[f(x,t)-g(t)^{2}\nabla_{x}\log p_{t}(x)\right]dt+g(t)\,d\overline{w},$ (11)
where $\nabla_{x}\log p_{t}(x)$ is the score function of the data distribution at time $t$, and $d\overline{w}$ is the reverse-time Wiener process. Since $\log p_{t}(x)$ is unknown, it is estimated using a score network $s_{\theta}(x,t)\approx\nabla_{x}\log p_{t}(x)$, trained using denoising score matching:
$\begin{split}\mathcal{L}(\theta)=&\;\mathbb{E}_{t\sim\mathcal{U}(0,T)}\lambda(t)\,\mathbb{E}_{x_{0},x_{t}}\big{[}\|s_{\theta}(x_{t},t)\\ &\;-\nabla_{x}\log p_{t}(x_{t}\mid x_{0})\|_{2}^{2}\Big{]}\end{split}$ (12)
where $\lambda(t)$ is a weighting function, $x_{0}\sim p_{\text{data}}$, and $x_{t}$ is the noisy version of $x_{0}$. Once trained, the model generates samples by solving the reverse SDE (e.g., using numerical solvers like Euler-Maruyama [108] or Predictor-Corrector schemes [109]), starting from Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ and integrating backward in time. These models have gained attention for their flexibility and effectiveness in capturing complex data distributions, making them well-suited for generating high-fidelity synthetic data across various domains. SOS [70] is the first score-based tabular data oversampling method, which redesigns the network to process tabular data, and through fine-tuning, the oversampling quality has been further improved. MissDiff [71] introduces a unified diffusion-based framework that incorporates a masked regression loss for denoising score matching during the training phase. This approach enables the model to effectively learn from data with missing values, addressing various missing data problems and improving the quality of synthetic data in the presence of incomplete information. To effectively process both categorical and numerical features simultaneously, CSDI-T [72] explores three techniques: one-hot encoding, analog bit encoding, and feature tagging. STaSy [73] leverages self-paced learning techniques and fine-tuning strategies to enhance the stability of denoising score-matching training, leading to improved sampling quality and greater diversity in synthetic tabular data. Forest-VP/Forest-Flow [74] employs XGBoost [110], leveraging its strength in tabular data prediction and classification. Importantly, XGBoost [110] can naturally handle missing data by learning the best split, allowing the model to be trained directly on incomplete datasets—an advantage over most generative models that require complete data. TABSYN [75] synthesizes tabular data by leveraging a diffusion model within a variational autoencoder (VAE) crafted latent space.
@@ III-C LLM-based Methods
Generating high-quality and realistic tabular data, particularly in specialized domains such as finance and healthcare, often requires domain-specific knowledge. For instance, generating synthetic heart disease records necessitates an understanding that the probability of heart disease increases with age, alongside general knowledge, such as the fact that a person’s age cannot be negative. However, existing generative models typically focus on learning the statistical distribution of raw tabular data without incorporating such knowledge. As a result, they may generate data that aligns with the learned distribution but violates logical or domain-specific constraints. To address this challenge, recent research explores leveraging the knowledge embedded in extensively pre-trained LLMs for synthetic tabular data generation. An overview of LLM-based models for synthetic tabular data generation is summarized in Table III. Depending on whether the LLM is fine-tuned, these methods can be broadly categorized into Prompt-based methods and Fine-Tuning methods as shown in Figure 5.
@@ III-C1 Prompt-based Methods
In-context learning (ICL) offers an innovative approach by enabling language models to perform tasks based solely on input-output examples, without parameter updates or fine-tuning. This capability, first observed in GPT-3 and subsequent LLMs [111], allows models to generalize to new tasks by embedding examples directly into the input prompt, which is often referred to as an "emergent ability" [112], this phenomenon has spurred significant research into understanding and enhancing ICL in models exceeding 100 billion parameters. Leveraging this capability, existing methods serialize the tabular dataset $D_{e}$ into a format compatible with LLM’s input. By carefully designing generation prompts, these methods guide the LLM to synthesize a synthetic tabular dataset $D_{s}$. EPIC [76] leverages balanced, grouped data samples and consistent formatting with unique feature mapping to guide LLMs in generating accurate synthetic data across all classes, even for imbalanced datasets. CLLM [16] leverages the prior knowledge of LLMs for data augmentation in the limited data scenarios, and introduces a principled curation mechanism, leveraging learning dynamics, coupled with confidence and uncertainty metrics, to obtain a high-quality dataset. LITO [77] proposes an oversampling framework for tabular data to guide the abilities of generative language models. By leveraging its conditional sampling capabilities, LITO [77] synthesizes minority samples by progressively masking the important features of the majority class samples and imputing them towards the minority distribution.
@@ III-C2 Fine-Tuning Methods
Directly prompting LLMs for synthetic tabular data generation may lead to inaccurate or misleading outputs that deviate from user instructions. To mitigate this issue, recent research has explored fine-tuning LLMs on tabular data, enabling them to better understand the structure, constraints, and relationships within tabular datasets. Fine-tuning helps improve the accuracy and coherence of synthetic tabular data, ensuring that the output aligns more closely with real-world data distributions and domain-specific knowledge. GReaT [79] fine-tunes LLM for sampling synthesized but highly realistic tabular data. In addition, GReaT can model the distribution of tabular data by adjusting any subset of features. TabMT [80] is a Bert-based mask transformer design used to generate synthetic tabular data, effectively addressing the unique challenges posed by heterogeneous data fields and capable of handling missing value problem. TAPTAP [81] utilizes tabular data pre-training to enhance the model’s tabular prediction ability. After pre-training with a large amount of real-world tabular data, TAPTAP [81] can generate high-quality tabular data to support various applications, including privacy protection, limited data scenarios, missing value imputation, and imbalanced classification. HARMONIC [82] constructs an instruction fine-tuning dataset based on the idea of the k-nearest neighbors algorithm to inspire LLMs to discover inter-row relationships. By fine-tuning, LLMs are trained to remember the format and connections of the data rather than the data itself, which reduces the risk of privacy leakage. AIGT [83] based on prompt enhancement, which utilizes metadata information, such as tabular descriptions and schemas, as prompts to generate a synthetic tabular dataset $D_{s}$ and propose long token partitioning algorithms that enable AIGT to model tabular data of any scale. DP-LLMTGen [17] for differentially private tabular data synthesis that leverages pre-trained large language models (LLMs) to model sensitive datasets using a two-stage fine-tuning procedure with a novel loss function specifically designed for tabular data. Subsequently, it generates synthetic data through sampling the fine-tuned LLMs. P-TA [87] proposes the use of Proximal Policy Optimization (PPO [113]) to apply GANs to guide LLMs to refine the probability distribution of tabular features and thereby generate more realistic data.
@@ IV Post Processing
The generative model introduced in Section III is theoretically capable of generating an infinite amount of tabular data. However, these synthetic samples often contain a number of unrealistic, incorrect, or even toxic instances. This issue arises from two main factors. First, the distribution of tabular data is highly complex, making its generation a traditionally challenging task. Second, non-LLM-based methods lack common sense and professional knowledge, leading to the generation of data that conforms to the statistical distribution but deviates from human knowledge. In LLM-based tabular generation methods, the "hallucination problem" exacerbates this issue, as LLMs inevitably produce flawed samples with incorrect labels. To address these challenges, several post-processing techniques for synthetic tabular data have been proposed. These methods can be broadly categorized into two groups: sample enhancement and label enhancement, as shown in Figure 6.
@@ IV-A Sample Enhancement
Sample enhancement methods aim to improve the quality and rationality of synthetic tabular data by modifying feature values in the synthetic samples or filtering the unreasonable sample to transform originally unreasonable or invalid samples into more realistic ones. They can be broadly categorized into two main approaches: sample filtering and sample correction.
@@ IV-A1 Sample Filtering
The core of sample filtering methods lies in identifying low-quality samples. Therefore, the key step is to design appropriate criteria based on learning dynamics [114, 115, 116], which may include factors such as confidence scores, impact functions, and generation capabilities, then samples are filtered according to these established criteria. For example, CLLM [16] analyzes learning dynamics to determine whether to sample or not by calculating two key metrics: confidence and uncertainty. Samples that are uncertain or produce inconsistent predictions can harm model performance, so they are filtered out. Similarly, LITO [77] uses self-authentication with rejection sampling. The language model itself can validate the correctness of synthetic samples by inferring labels and filtering out inaccurate or inconsistent ones, improving the reliability of the data without external validation.
@@ IV-A2 Sample Correction
Sample correction incorporates neuro-symbolic principles into the training process of tabular data generation models by introducing human knowledge in the form of constraints or rules. Such as, LL [117] addresses the issue of aligning synthetic samples with background knowledge by adding an additional layer to restrict the output space to comply with a set of linear inequalities, ensuring consistency between synthetic data and available domain knowledge. However, this approach is limited to cases where the knowledge can be accurately expressed using linear inequalities, whose representational capacity is inherently constrained. To overcome this limitation, DRL [118] extends the approach by representing background knowledge using Quantifier-Free Linear Real Arithmetic (QFLRA) formulas, allowing the model to enforce constraints in non-convex and even disconnected spaces.
@@ IV-B Label Enhancement
Label enhancement methods aim to correct potential annotation errors in synthetic samples. The most straightforward approach is manual re-annotation of mislabeled data. Wang et al. [28] propose actively selecting samples with the lowest confidence scores for human re-labeling, while Pangakis et al. [119] and Liu et al. [120] further emphasize the importance of manual review in ensuring data quality. However, despite its effectiveness, manual re-labeling incurs significant costs and is often impractical for large-scale deployment. To address these limitations, most synthetic tabular data generation methods adopt the proxy model approach, which leverages automated models to refine and correct labels. TAPTAP [81] and AIGT [83] use generative models to generate tabular features and then use the backbone discriminative model trained on the original data to generate pseudo labels for the synthetic data. Pred-LLM [84] generates tabular features through pre-training LLM, while further prompting LLM to generate its labels.
@@ V Evaluation
Kaggle Churn CH 10000 11(7/4) None Binclass Finance
@@
The performance of tabular data generative models is typically evaluated from multiple, complementary perspectives. In this section, we categorize the evaluation of synthetic tabular data generation into two dimensions: Data Availability and Privacy Protection. Data availability refers to the effectiveness of the synthetic data in supporting downstream machine learning tasks, including learning efficiency, fidelity to the original data distribution, and alignment with human domain knowledge. Privacy Protection, on the other hand, measures the extent to which the synthetic data avoids disclosing sensitive information that could potentially be used to re-identify user information from the original dataset.
@@ V-B1 ML Efficiency
Machine Learning (ML) utility is a key criterion for evaluating the effectiveness of synthetic tabular data, as it reflects the extent to which the synthetic data can support downstream learning tasks. One widely adopted evaluation protocol for this purpose is the Training on Synthetic and Testing on Real (TSTR) scheme [29]. In this framework, tabular data is first divided into a training set and a test set. A generative model is then trained on the training data and used to produce a synthetic tabular dataset $D_{s}$ of the same size. Subsequently, two machine learning models are trained separately: one on the real training data and the other on the synthetic data. Both models are then evaluated using the same real test set. The core assumption behind this scheme is that if the synthetic data accurately captures the underlying patterns and distributions of the original data, then a model trained on the synthetic data should generalize well to real-world data. Therefore, if the performance of the model trained on synthetic data is comparable to that of the model trained on real data (e.g., in terms of accuracy, F1 score, AUC, and RMSE, etc.), the synthetic tabular dataset $D_{s}$ is deemed to possess high machine learning utility. notably $\text{AUC}=\int_{0}^{1}\text{ROC}(t)\,d(t)$ is commonly used for evaluating categorical data, while $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ is commonly used for regression datasets.
@@ V-B2 Fidelity
Fidelity measures how well synthetic data preserves the statistical properties of the tabular data. Although recent studies [30, 31] suggest that high fidelity does not always imply strong downstream task performance, it remains a crucial aspect of evaluating synthetic data quality.
@@
  • Column-wise: This metrics compares individual features between real and synthetic data. Common metrics include the Kolmogorov-Smirnov Test (KST) [121] for numerical features. Given two numerical distributions, $p_{r}(x)$ (representing real data) and $p_{s}(x)$ (representing synthetic data), KST quantifies the distance between these distributions by calculating the maximum discrepancy between their corresponding Cumulative Distribution Functions (CDFs):
    $\text{KST}=\sup_{x}\left|F_{r}(x)-F_{s}(x)\right|,$ (13)
    where $F_{r}(x)$ and $F_{s}(x)$ are the CDFs of $p_{r}(x)$ and $p_{s}(x)$, respectively:
[…]
    For categorical data, the Total Variation Distance (TVD) [122] is commonly used. TVD computes the frequency of each category and expresses it as a probability. The TVD score represents the average difference between the probabilities of each category:
    $\text{TVD}=\frac{1}{2}\sum_{\omega\in\Omega}\left|R(\omega)-S(\omega)\right|,$ (15)
    where $\omega$ represents all possible categories in a given column $\Omega$, and $R(\cdot)$ and $S(\cdot)$ denote the real and synthetic frequencies for these categories, respectively.
  • Pair-wise: Assesses how well dependencies between feature pairs are preserved using metrics such as Pearson Correlation Coefficient Matrices (DPCM) (for numerical features), The difference between Contingency Similarity Matrices among categorical features (DCSM) (for categorical features), and The difference between Contingency Similarity Matrices (for numerical and categorical features).